Suppose in Flower Land people grow roses, and you have a demand and supply curve for roses, where P is the price for roses and Q is the quantity of roses: P=120-8Q P=10+2Q Please find the equilibrium price and quantity for roses. Please graph supply and demand curves and show the equilibrium price and quantity demanded on the graph. Please also label the axes, intercepts, and curves.
Answer #1

Given, demand: P = 120-8Q and supply: P = 10+2Q, equilibrium is atrained at a point where demand = supply.

Thus, 15- P/8 = 5+P/5

which gives P = 32 and Q = 11

Date 1o- ci.